OpenStudy (anonymous):

I toss a coin 4 times. Find the chance of getting: a. the sequence HTHT b. 2 heads c. more heads than tails Answers please :)

OpenStudy (anonymous):

a. H: Head T: Tail Sequence: Head, Tail, Head, Tail.

OpenStudy (anonymous):

1) P(HTHT)=0.5^4=0.0625---->6.25% (1 out of 16) 2)P(2 Heads) as it is a binomial distribution:\[P(k=2)=\left(\begin{matrix}4 \\ 2\end{matrix}\right)0.5^2·0.5^{4-2}=6·0.0625=0.375\], that is a 37.5% or 6 chances out of 16 3)More heads than tails is that you get either zero or one tail:\[P(k<2)=\left(\begin{matrix}4 \\ 0\end{matrix}\right)0.5^0·0.5^4+\left(\begin{matrix}4\\ 1\end{matrix}\right)0.5^1·0.5^{(4-1)}=0.3125\] that is, 5 chances out of 16 A manual approach to solve the problem: HHHH case (c) HHHT case (c) HHTH case (c) HHTT case (b) HTHH case (c) HTHT case (b) HTTH case (b) HTTT THHH case (c) THHT case (b) THTH case (a), case (b) THTT TTHH case (b) TTHT TTTH TTTT Case(a) 1 out of 16 Case (b) 6 out of 16 Case (c) 5 out of 16

OpenStudy (anonymous):

Many thanks! :)

OpenStudy (anonymous):

The important thing is not the solution, is that you get the idea.

OpenStudy (anonymous):

trust me i got the idea :)

OpenStudy (anonymous):

happy to know .That was the goal

OpenStudy (anonymous):

Solution Problem 1 A,B & C and 4B http://0e86bde3.linkbucks.com

OpenStudy (anonymous):

Solution PROBLEM SET 3 ANSWERS ==> HTTP://F96DBEB2.LINKBUCKS.COM PROBLEM SET 4 ANSWERS ==> HTTP://07891A62.LINKBUCKS.COM